# Formation of Bound States in Quintessence Alternative Theories

### v1

@@ V Formation of Bound States
We now proceed with the calculation of the masses of the bound states, along with their imaginary parts. The latter signal the instability of these states, since they may tunnel through the barrier on the left and end up inside the horizon. To find the real part of the masses we follow Grain:2007gn and consider the equation:
[…]
@@
To determine the imaginary parts of these states, interpreted as bandwidths, we fix the value of $\omega$ that we found in the previous step (call it $\tilde{\omega}$ say), and we calculate the bandwidth via:
[…]
@@
$\frac{d^{2}(p^{2})}{dr_{*}^{2}}=f^{2}(r)\frac{d^{2}(p^{2})}{dr^{2}}+f(r)f^{\prime}(r)\frac{d(p^{2})}{dr},\ p^{2}=\tilde{\omega}^{2}-V_{0}.$ (35)
Notice that $r_{1}$ and $r_{2}$ are the two points common to the horizontal line at $\tilde{\omega}^{2}$ and the potential $V_{0},$ which lie nearest to the origin, where the $\tilde{\omega}$ horizontal line cuts the potential curve, while $r_{3}$ is the most distant cut. In addition, $r_{max}$ is the point, where the potential attains its maximal value. After we are done with these calculations, we have the mass eigenvalues in the form $\tilde{\omega}-i\gamma.$
We now report the mass eigenvalues and their bandwidths for various values of $w$ and $a.$ In Fig. 6, left panel, we depict the dependence of the masses on $w$ for $Q=0.00,\ M=0.05,\ a=0.8$ and $l=0.$ It appears that there exist no bound states for $|w|$ smaller than some value (about 0.55), while the bandwidths (denoted by error bars) increase as we approach this value. The masses do not change much, unless near the value $w=-1.$ In the right panel we show the corresponding results for $Q=0.00,\ M=0.05,\ w=-0.6$ and $l=0.$ There is an almost linear dependence of the masses on $a,$ while the bandwidths are more or less constant.
Figure 6: Left panel: Masses versus $w$ for $l=0.$ Right panel: Masses versus $a$ for $l=0.$
In Fig. 7, left panel, we depict the dependence of the masses on $w$ for $Q=0.00,\ M=0.05,\ a=0.8$ and $l=1.$ This larger angular momentum allows for eigenvalues for values of $w$ smaller than before. In addition the bandwidths are very much smaller and this is due to the large barrier on the left due to the angular momentum (compare Fig. 5); this barrier does not allow the particles go into the horizon easily, thus their imaginary parts are reduced to the size of the symbols. The masses are larger than before and the qualitative behaviour versus $w$ is the same. In the right panel we depict the corresponding results for $Q=0.00,\ M=0.05,\ w=-0.6$ and $l=1.$ The dependence of the masses on $a,$ is once more linear.
Figure 7: Left panel: Masses versus $w$ for $l=1.$ Right panel: Masses versus $a$ for $l=1.$
@@ VI Conclusions
We have studied black hole solutions in the presence of quintessence. In particular we concentrated in the bound states of scalar excitations outside the horizon. It has been shown that the parameters of the quintessence influence significantly the metric function and the Hawking temperature. It has turned out that large black holes radiate less for large $|w|,$ while small black holes radiate more. Thus large black holes seem to be longer lived, compared to the ones with no quintessence for large $|w|.$
The Regge-Wheeler potentials depend on the charge $Q$ of the black hole: when $Q$ gets large enough, there is no left barrier, so the scalar particles will meet no obstacle to dive into the black hole. This has nothing to do with the quintessence. The presence of quintessence confines them to smaller values in general, although the qualitative picture of the $Q$ dependence does not change. For small enough $Q$ thw presence of the quintessence may induce qualitative changes: the potential for $w=-0.40$ does not support scalar bound states, but increasing $|w|$ may create a potential with a barrier on the left, which can support bound states.
A barrier on the left may also be enhanced by increasing the angular momentum quantum number $l.$ This suggests that the scalar particles with large angular momentum are longer lived than tho ones with smaller $l.$ The study of the potentials suggest that $|w|$ should be larger than some bound, if bound states are to be supported. On the other hand the masses decrease for increasing parameter $a$ almost linearly. Also, the states with large angular momenta $l$ have larger masses and smaller bandwidths than the ones with small $l.$

### v2

@@ V Calculation of Bound States
We now proceed with the calculation of the masses of the bound states, along with their imaginary parts. The latter signal the instability of these states, since the latter may tunnel through the barrier on the left and end up inside the horizon. To find the real part of the masses we follow Grain:2007gn and consider the equation:
$\int_{r_{2}}^{r_{3}}\frac{\sqrt{\omega^{2}-V_{0}}}{f(r)}dr=\left(n+\frac{1}{2}\right)\pi,n=0,\ 1,\ 2,\ \dots.$ (31)
and we tune $\omega,$ so that the equation is satisfied for some integer $n.$ The states with $n=0$ are the fundamental massive modes, while the ones with n=1, 2,n=1,\ 2,\ $\dots$ are the overtones. In general a small number of overtones exists, since the energy levels eventually get higher than the left barrier. We also consider various values of the angular momentum quantum number $l;$ we will show the results for $l=0$ and $l=1.$
@@
To determine the imaginary parts of these states, interpreted as bandwidths, we fix the real part of $\omega,$ that we found in the previous step (call it $m$), and we calculate the bandwidth via:
$\gamma=m\frac{\exp\left(-2\int_{r_{1}}^{r_{2}}\frac{\sqrt{V_{0}-m^{2}}}{f(r)}dr\right)}{1+\exp\left(-\frac{2\pi p^{2}(r_{max})}{\sqrt{2\frac{d^{2}p^{2}(r_{max})}{dr_{*}^{2}}}}\right)}.$ (32)
@@
$\frac{d^{2}(p^{2})}{dr_{*}^{2}}=f^{2}(r)\frac{d^{2}(p^{2})}{dr^{2}}+f(r)f^{\prime}(r)\frac{d(p^{2})}{dr},\ p^{2}=m^{2}-V_{0}.$ (33)
Notice that $r_{1}$ and $r_{2}$ are the two points common to the horizontal line at $m^{2}$ and the potential $V_{0},$ which lie nearest to the origin, where the $m$ horizontal line cuts the potential curve, while $r_{3}$ is the most distant cut. In addition, $r_{max}$ is the point, where the potential attains its maximal value. After we are done with these calculations, we have the mass eigenvalues in the form $m-i\gamma.$
We now report the mass eigenvalues and their bandwidths for various values of $w$ and $a.$ In Fig. 6, left panel, we depict the dependence of the masses on $w$ for Q=0.00,M=0.05,a=0.8Q=0.00,\ M=0.05,\ a=0.8 and $l=0.$ It appears that there exist no bound states for $|w|$ smaller than some value (about 0.55), while the bandwidths (denoted by error bars) increase as we approach this value. The masses do not change much, unless near the value w=1.w=-1. In the right panel we show the corresponding results for Q=0.00,M=0.05,w=0.6Q=0.00,\ M=0.05,\ w=-0.6 and $l=0$ versus $a.$ There is an almost linear dependence of the masses on $a,$ while the bandwidths are more or less constant. This behaviour must be correlated with the changes in the potentials, brought about by the changes in $w.$ In Fig. 4, left panel, it is seen that, as $|w|$ decreases, the potential gradually loses its left barrier. This is why we are not able to find masses for $|w|$ smaller than some value. In addition, in Fig. 5, right panel, it is seen that the potentials get larger values for small values of $a,$ so one would expect larger values for the masses in this region, which is actually observed in Fig. 6.
Figure 6: Left panel: Masses versus $w$ for Q=0.00,M=0.05,a=0.8Q=0.00,\ M=0.05,\ a=0.8 and $l=0.$ Right panel: Masses versus $a$ for Q=0.00,M=0.05,w=0.6Q=0.00,\ M=0.05,\ w=-0.6 and $l=0.$
In Fig. 7, left panel, we depict the dependence of the masses on $w$ for Q=0.00,M=0.05,a=0.8Q=0.00,\ M=0.05,\ a=0.8 and $l=1.$ This larger angular momentum allows for eigenvalues at values of $|w|$ smaller than before. This is due to the fact that the decrease in the height of the barrier on the left, due to the decrease of $|w|,$ is partially compensated by the increase, due to the larger angular momentum. In addition the bandwidths are very much smaller and this is due to the large barrier on the left due to the angular momentum (compare Fig. 5); this barrier does not allow the particles go into the horizon easily, thus their imaginary parts are reduced to the size of the symbols. The masses are larger than before and the qualitative behaviour versus $w$ is the same. In the right panel we depict the corresponding results for Q=0.00,M=0.05,w=0.6Q=0.00,\ M=0.05,\ w=-0.6 and $l=1$ versus $a.$ The dependence of the masses on $a$ is, once more, linear.
Figure 7: Left panel: Masses versus $w$ for Q=0.00,M=0.05,a=0.8Q=0.00,\ M=0.05,\ a=0.8 and $l=1.$ Right panel: Masses versus $a$ for Q=0.00,M=0.05,w=0.6Q=0.00,\ M=0.05,\ w=-0.6 and $l=1.$
@@ VI Conclusions
We have studied black hole solutions in the presence of quintessence. In particular we concentrated in the bound states of scalar excitations outside the horizon. It has been shown that the parameters of the quintessence influence significantly the metric function and the Hawking temperature. It has turned out (Fig. 2) that, for w=0.40,w=-0.40, large black holes $(r\gtrsim 0.6)$ radiate more than small ones, while, as $|w|$ increases, the behaviour changes: for w=1.00,w=-1.00, small black holes radiate more than big ones. Thus large black holes with large $|w|$ seem to be longer lived, compared to large black holes with small $|w|.$ The situation is reversed for small black holes. In the long run, one may expect to find mainly large black holes with large $|w|,$ or small black holes with small $|w|,$ if other factors (e.g. angular momentum or charge) do not change. In addition, black holes with large $a$ have smaller temperatures and radiate less than the ones with small (or not at all) $a.$
The Regge-Wheeler potentials depend on the charge $Q$ of the black hole: when $Q$ gets large enough, there is no left barrier, so the scalar particles will meet no obstacle to dive into the black hole (Figs. 3 and 4). This has nothing to do with the quintessence. The presence of quintessence enhances the left barrier to smaller values. For $Q$ values modest enough, the presence of the quintessence may induce qualitative changes: the potential for w=0.40w=-0.40 does not support scalar bound states, but larger $|w|$ may create a potential with a barrier on the left, which can support bound states. The barrier on the left may be enhanced by increasing the angular momentum quantum number $l$ (Fig. 5, left panel). This suggests that the scalar particles with large angular momentum are longer lived than the ones with smaller $l.$
The study of the potentials suggests that $|w|$ should be bigger than some bound, if bound states are to be supported, as depicted in Fig. 6, left panel. The presence of quintessence reduces the values of the potentials (Fig. 5, right panel), so we expect that the masses of the possible bound states will decrease with increasing $a.$ The decrease for increasing parameter $a$ is almost linear (right panel of the same figure). Also, the states with large angular momenta $l$ have larger masses and smaller bandwidths than the ones with small $l.$ The relevant results are to be found in Fig. 7. In short, bound states emerge at $a=0,$ while if $a\neq 0,$ bound states are only present for large enough $|w|$ and/or $l.$
As it is well known Chiba:1999wt , the observational evidence for a cosmological constant, which is responsible for the expansion of the Universe, gives the bounds 1<w<1/3.-1<w<-1/3. In accord with this, we have found that bound states may be formed for large absolute values of the quintessence parameter $|w|,$ contributing in this way to the expansion of the Universe.
In our study we have not considered extra matter surrounding the quintessence black hole. If this matter distribution does not back-react on the background metric, the no-hair theorems do not allow to influence the structure of the background black hole. However, if we perturb this matter distribution calculating the QNMs, an observer outside the black hole will detect the properties of the quintessence matter. We leave this study for future work.